Write the equation 2x^2-16x-y+24=0 in vertex form.
Find the..
(a) vertex
(b) axis of symmetry
(c) focus
(d) directrix oft the parabola

thank you.

Answer by stanbon(57284)

Write the equation 2x^2-16x-y+24=0 in vertex form.
Find the..
(a) vertex
(b) axis of symmetry
(c) focus
(d) directrix oft the parabola
------------------------------
2x^2-16x-y+24=0
2x^2-16x = y-24
2(x^2-8x+16) = y-24+2*16
2(x-4)^2 = y + 8
(x-4)^2 = (1/2)(y+8)
---------------------
(a) vertex:.................. (4,-8)
(b) axis of symmetry:........ x = 4
(c) focus....................
4p = 1/2 ; p = 1/8
So, focus is (4,-8+(1/8)) or (4,-63/8)
(d) directrix of the parabola.. y = -8-(1/8) = -65/8
=============================================
Cheers,
Stan H.